three capacitors c1 = 2uf ,c2= 4uf, c3= 4uf, are connected in series and parallel, find the total capacitance for 3 capacitors


1
Expert's answer
2022-03-02T10:15:13-0500

When capacitors are connected in parallel, the total capacitance is equal to the sum of the capacitance of each of the capacitors. Hence Ctotal=C1+C2+C3.

= 2+4+4=10uf